OpenStudy (anonymous):

Can anybody help me with this equation please? y = 1E-09x3 - 6E-06x2 + 0.0089x + 1.2673 Basically I want to rewrite it simply, without letters inside, but with mathematical symbols and abbreviations (e.g "exp" etc). Any ideas please?

OpenStudy (unklerhaukus):

y = 1E-09x3 - 6E-06x2 + 0.0089x + 1.2673 \[y = 1\times10^{-9}x^3 - 6\times10^{-6}x^2 + 0.0089x + 1.2673\\ \quad=\frac{x^3}{10^9}-6\frac{x^2}{10^6} + 8.9\frac{x}{10^3} + 1.2673\\ \quad=\left(\frac{x}{10^3}\right)^3-6\Big(\frac{x}{10^3}\Big)^2 + 8.9\big(\frac{x}{10^3}\big) + 1.2673\\\]

OpenStudy (anonymous):

Thank you very much for your great help! If I face something else I will come back. Mathematics for ever!

OpenStudy (unklerhaukus):

ok